you are making a mosic desig on a square table top. Yuo have already covered half of the table top with 150 1 inch square tile pieces.
what are the dimensions of the table top?
what is the measure of the diagonal from one corner to the opposite corner of the table top?

Answer by stanbon(75887)

you are making a mosic desig on a square table top. Yuo have already covered half of the table top with 150 1 inch square tile pieces.
what are the dimensions of the table top?
---
Total table area = 300 sq in
----
since it's a square the dimensions are equal
width = length = sqrt(300) = sqrt(100*3) = 10sqrt(3) inches
------------------------------------
what is the measure of the diagonal from one corner to the opposite corner of the table top?
Use Pythagoras to get:
diag = sqrt[(sqrt(300))^2 + (sqrt(300))^2] = sqrt(600)= 10sqrt(6)
=========================
Cheers,
Stan H.